OpenStudy (anonymous):

Given the function f described by f(y)=4y^2+9y-3 find the function values f(2)

OpenStudy (anonymous):

substitute 2 for y in the equation and calculate the value... f(y)=4y^2+9y-3 so f(2) = 4(2)^2 + 9(2) - 3 = 16 + 18 - 3 = 31